# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 3. $(3)$
(and similarly with $l^{*}_{+}$, $_{+}, m+m^{\prime}_{+}$). So replacing $O$ with the subgraph that remains upon removal of the two endpoints works. ∎
The general case of Lemma B.1 will be reduced to the lemma above through the following lifting argument.
Lemma B.3.
Let $L$ be a finite linear graph, $l^{+},l^{-}$ be two nodes and $^{+}, tt^{-}$ be two $L$-types in the sense of Definition 5.5 starting with $^{+}, ll^{-}$, respectively, that is, $t^{-}$ is of the form $\{\{l^{-}_{i}:\ i<k\}:\ k=1,2,\ldots,|L|\}$, an increasing chain of connected subgraphs of $L$ with $l_{0}=l^{-}$ and $\{l^{-}_{i}:\ i<|L|\}=L$.
Then there exists a finite linear graph $L^{*}$, and epimorphism $h\in{\rm Epi}(L^{*},L)$, with endpoints $l^{*}_{-},l^{*}_{+}$ of $L^{*}$ satisfying ${\rm tp}^{l^{*}_{-},L^{*}}(h)=t^{-}$, ${\rm tp}^{l^{*}_{+},L^{*}}(h)=t^{+}$, such that whenever $M$ is a finite linear graph with endpoints $m-,m_{+}\in M$ and $f\in{\rm Epi}(M,L)$ that satisfies ${\rm tp}^{m_{+},M}(f)=t^{+}$, ${\rm tp}^{m_{-},M}(f)=t^{-}$, then there exists a finite linear graph $M^{*}$, endpoints $m^{*}_{+},m^{*}_{-}\in M^{*}$, epimorphisms $g\in{\rm Epi(M^{*},M)}$ $f^{*}\in{\rm Epi}(M^{*},L^{*})$ with
  • $f\circ g=h\circ f^{*}$,
  • $g(m^{*}_{+})=m_{+}$,
  • $g(m^{*})=m_{-}$,
  • $f^{*}(m^{*}_{+})=l^{*}_{+}$,
  • $f^{*}(m^{*}_{-})=l^{*}_{-}$,
(in particular both $f^{*}$ and $g$ map endpoints to endpoints).
Proof..
The lifting will be done in two steps, handling only one endpoint at a time.
Claim B.4.
Assume that $L$, $^{+}, tt^{-}$ are as in the lemma. Then there exists a finite linear graph $L^{*}$, an epimorphism $h\in{\rm Epi}(L^{*},L)$, an endpoint $l^{*}_{-}$, and an $L^{*}$-type
$(t^{*})^{+}=\{\{(l^{*})^{+}_{j}:j<k\}:\ k\leq|L^{*}|\},$
such that $h((t^{*})^{+})=t^{+}$, moreover, if $l^{+}_{0}$ is an endpoint of $L$, then so is $(l^{*})^{+}_{0}$, and $(l^{*})^{+}_{0}\neq l^{*}_{-}$.
Moreover, whenever $M$ is a finite linear graph with endpoints $m-,m_{+}\in M$ and $f\in{\rm Epi}(M,L)$ that satisfies ${\rm tp}^{m_{+},M}(f)=t^{+}$, ${\rm tp}^{m_{-},M}(f)=t^{-}$, then there exists a finite linear graph $M^{*}$, endpoints $m^{*}_{+},m^{*}_{-}\in M^{*}$, epimorphisms $g\in{\rm Epi(M^{*},M)}$ $f^{*}\in{\rm Epi}(M^{*},L^{*})$ with
  • $f\circ g=h\circ f^{*}$,
  • $g(m^{*}_{+})=m_{+}$,
  • $g(m^{*})=m_{-}$,
  • ${\rm tp}^{m^{*}_{+},M^{*}}(f^{*})=(t^{*})^{+}$,
  • $f^{*}(m^{*}_{-})=l^{*}_{-}$ (in particular, ${\rm tp}^{m^{*}_{-}}(f^{*})$ is the unique $L^{*}$-type starting with $l^{*}_{-}$).
First we construct $L^{*}$ and $h:L^{*}\to L$. The linear graph $L^{*}$ will be the disjoint union of the linear graphs $P$, $Q$ and $S$ with
  1. $(\nabla_{P})$
    $P=\{p_{0},p_{1},\ldots,p_{|P|-1}\}$,
    1. $(\nabla_{P})_{1}$
      $p_{i}$ is connected with $p_{j}$ iff $|i-j|\leq 1$,
    2. $(\nabla_{P})_{2}$
      $p_{0}$ will be an endpoint of $L^{*}$,
    3. $(\nabla_{P})_{3}$
      $h(p_{0})=l_{-}$,
    4. $(\nabla_{P})_{4}$
      ${\rm tp}^{p,P}(h)=t^{-}$, in particular, $h{\restriction}P$ is surjective,
    5. $(\nabla_{P})_{5}$
      $h(p_{|P|-1})=l^{-}_{|L|-1}$, is an endpoint of $L$,
  1. $(\nabla_{Q})$
    $Q=\{q_{0},q_{1},\ldots,q_{|L|-1}\}$,
    1. $(\nabla_{Q})_{1}$
      $q_{i}$ is connected with $q_{j}$ iff $|i-j|\leq 1$,
    2. $(\nabla_{Q})_{2}$
      $q_{0}$ will be an endpoint of $L^{*}$,
    3. $(\nabla_{Q})_{3}$
      $h{\restriction}Q:Q\to L$ is an isomorphism (so $h(q_{0})$ is an endpoint of $L$),
    4. $(\nabla_{Q})_{4}$
      $h(q_{|L|-1})=l^{+}_{|L|-1}$ is the other endpoint of $L$,
  1. $(\nabla_{S})$
    $S=\{s_{0},s_{1},\ldots,s_{|S|-1}\}$ is a linear graph
    1. $(\nabla_{S})_{1}$
      $s_{i}$, $s_{j}$ is connected iff $|i-j|\leq 1$,
    2. $(\nabla_{S})_{2}$
      $s_{0}$ is connected with $p_{|P|-1}$, and $h(s_{0})=h(p_{|P|-1})$,
    3. $(\nabla_{S})_{3}$
      $s_{|S|-1}$ is connected with $q_{|L|-1}$, and $h(s_{|S|-1})=h(q_{|L|-1})$,
    4. $(\nabla_{S})_{4}$
      $h{\restriction}S:S\to L$ is surjective,
    5. $(\nabla_{S})_{5}$
      if $h(p_{|P|-1})=h(q_{|Q|-1})$, then |S|=2|L|1|S|=2|L|-1,
    6. $(\nabla_{S})_{6}$
      if $h(p_{|P|-1})\neq h(q_{|Q|-1})$ (so they are necessarily opposite endpoints of $L$), then $S$ is the linear graph with $|S|=|L|$,
Note that $h(s_{|S|-1})$, $h(s_{0})$ are endpoints, and so it follows from the surjectivity of $h{\restriction}S$ ($(\nabla_{S})$$(\nabla_{S})_{4}$) together with $(\nabla_{S})$$(\nabla_{S})_{5}$, $(\nabla_{S})$$(\nabla_{S})_{6}$ that $h(s_{|S|-1})$, $h(s_{0})$ fully determine $h{\restriction}S$.
We construct $P$ as follows. The linear graph $P$ will be the disjoint union of the intervals $P_{0},P_{1},\ldots,P_{k-1}$ for some $k$ (where $P_{i}$ is connected with $P_{i-1}$, and $P_{i+1}$).
We recall that $l_{0}=l_{-}$, and $\{l^{-}_{i}:\ i<k\}$ ($k=1,2,\ldots,|L|$) is an increasing chain of connected subgraphs (intervals) of $L$, coding the type $t^{-}$. We define the sequence $i_{0},i_{1},\ldots i_{j_{0}}$ (as well as $j_{0}$) by recursion as follows. First, set $i_{0}=0$, and assuming that $i_{j}$ is defined we let $i_{j+1}$ be minimal such that $i_{j+1}>i_{j}$ and $\neg(l^{-}_{i_{j+1}-1}Rl^{-}_{i_{j+1}})$, if such $i_{j+1}$ exists. If this cannot be continued, and we have defined $i_{0},i_{1},\ldots,i_{j}$, then define $i_{j+1}=|L|$, and let $j_{0}=j+1$), and we stipulate that $P$ will be the disjoint union of $P_{j}$’s, where $j<j_{0}$. (Note that $j_{0}\geq 2$, otherwise we are in the case of Lemma B.2).
Let $P_{0}=\{p_{0},p_{1},\ldots,p_{i_{1}-1}\}$ be a copy of $\{l^{-}_{0},l^{-}_{1},\ldots,l^{-}_{i_{1}-1}\}$, so it has endpoints $l^{-}_{0}=l_{-}$, and $l^{-}_{i_{1}-1}$, and define $h(p_{j})=l^{-}_{j}$ for $j<i_{1}$, i.e.
  1. $(\bigstar)_{0}$
    $h{\restriction}P_{0}$ is one-to-one, maps $P_{0}$ onto the interval between $l^{-}_{0}=h(p_{0})$ and $l^{-}_{i_{1}-1}=h(p_{i_{1}-1})$, and for future reference we note the following:
    $h[P_{0}]=\{l^{-}_{0},l^{-}_{1},\ldots,l^{-}_{i_{1}-1}\}.$
For $j<j_{0}$ we let $P_{j}=\{p_{|P_{<j}|},p_{|P_{<j}|+1},\ldots,p_{|P_{<j}|+|P_{j}|-1}\}$ (where $P_{<j}$ stands for $\bigcup_{k<j}P_{k}$), and define $h{\restriction}P_{j}$ in such a way that for each $j<j_{0}$ (with $j\geq 1$) the following hold.
  1. $(\bigstar)_{j}$
    $h{\restriction}(P_{j}\cup\{p_{|P_{<j}|-1}\}$ is one-to-one, mapping $P_{j}\cup\{p_{|P_{<j}|-1}\}$ onto the interval between $l^{-}_{i_{j}-1}=h(p_{|P_{<j}|-1})$ and $l^{-}_{i_{j+1}-1}=h(p_{|P_{<j}|+|P_{j}|-1})$, and this interval can be written as
    $h[P_{j}\cup(\{p_{|P_{<j}|-1}\}]=\{l^{-}_{0},l^{-}_{1},\ldots,l^{-}_{i_{j+1}-1}\}.$
Note that we defined $i_{j_{0}}$ to be $|L|$, so
  1. $(\bigstar)_{j_{0}}$
    $l^{-}_{i_{j_{0}}-1}=l^{-}_{|L|-1}=h(p_{|P|-1})$ is necessarily an endpoint of $L$.
Moreover, $l^{-}_{(i_{j_{0}-1})-1}$ must be the other endpoint of $L$. It is not difficult to see that ${\rm tp}^{p_{0},P}(h)=t^{-}$. This finishes the construction of $P$ and $h{\restriction}P$.
It is not difficult to see that there is a unique $Q$ and $h{\restriction}Q$ satisfying $(\nabla_{Q})$$(\nabla_{Q})_{1}$-$(\nabla_{Q})$$(\nabla_{Q})_{4}$.
So by the fact that $h(q_{|L|-1})=l^{+}_{|L|-1}$ is an endpoint of $L$ ($(\nabla_{Q})$$(\nabla_{Q})_{4}$) and $(\bigstar)_{j_{0}}$ the interval $S$, lying in between $P$ and $Q$ can be defined as required in, $(\nabla_{S})$$(\nabla_{S})_{4}$, $(\nabla_{S})$$(\nabla_{S})_{5}$, $(\nabla_{S})$$(\nabla_{S})_{6}$, i.e.
  1. $(\bigstar)_{S}$
    if $h(s_{0})=h(p_{|P|-1})$ and $h(q_{|L|-1})=h(s_{|S|-1})$ are opposite endpoints of $L$, then $|S|=L$, and $h{\restriction}S$ is a bijection, or else, if $h(p_{|P|-1})=h(q_{|L|-1})=h(s_{|S|-1})$, then |S|=2|L|1|S|=2|L|-1 and
    $h{\restriction}\{s_{0},s_{1},\ldots,s_{|L|-1}\}\;\hbox{ and }\;h{\restriction}\{s_{|L|-1},s_{|L|},\ldots,s_{2|L|-1}\}$
    are both bijections with $L$.
Now, having constructed $L^{*}=P\cup S\cup Q$ and $h$, fix $M$, $f\in{\rm Epi}(M,L)$, we need to construct $M^{*}$, $g\in{\rm Epi}(M^{*},M)$ and to lift up $f$ to some $f^{*}:M^{*}\to L^{*}$ with $h\circ f^{*}=f\circ g$.
$M^{*}$ will be defined to be the disjoint union of the subgraphs $U$, $V$ and $W$, where $f^{*}[U]=P$, $f^{*}[V]=Q$, $f^{*}[W]=S$. We begin with constructing $U$, and $f^{*}{\restriction}U$. Using $j_{0}$ from the construction of $P$, and recalling that ${\rm tp}^{m_{+},M}(f)=t^{+}$ we define the sequence $k_{0},k_{1},\ldots,k_{j_{0}}$ as follows. First we fix
(B.3) $\begin{array}[]{l}\textrm{an enumeration }\{m_{0},m_{1},\ldots,m_{|M|-1}\}\textrm{ of }M\\ \textrm{with }m_{0}=m_{+},\ m_{|M|-1}=m_{-}\textrm{ and }m_{j}Rm_{j+1}.\end{array}$
Let $k_{0}=0$, and let $k_{j}$ be minimal $k$ such that $f(m_{k})=l^{-}_{i_{j}-1}$. Note that ${\rm tp}^{m_{+},M}(f)=t^{-}$ implies that $k_{j}<k_{j+1}$. Now we note that $f$ maps the interval $M_{0}:=\{m_{0}=m_{k_{0}},m_{1},\ldots,m_{k_{1}}\}$ onto the interval between $f(m_{0})=l^{-}_{0}$ and $f(m_{k_{1}})=l^{-}_{i_{1}-1}$, and similarly $h$ maps $P_{0}$ onto this interval, endpoints $p_{0}$ and $p_{l_{i_{1}-1}}$ being mapped to $l^{-}_{0}$ and $l^{-}_{{i_{1}}-1}$ by $(\bigstar)_{j}$. Therefore, we can invoke Lemma B.2, and obtain
  1. $(\hexstar)_{0}$
    $U_{0}=\{u_{0},\ldots,u_{|U_{0}|-1}\}$, and $f^{*}{\restriction}U_{0}\in{\rm Epi}(U_{0},P_{0})$, $g{\restriction}U_{0}\in{\rm Epi}(U_{0},M_{0})$ with
    • $h\circ f^{*}{\restriction}U_{0}=f\circ g{\restriction}U_{0}$,
    • $f^{*}(u_{0})=p_{0}$,
    • $f^{*}(u_{|U_{0}|-1})=p_{|P_{0}|-1}(=p_{i_{0}-1})$,
    • $g(u_{0})=m_{0}(=m)$,
    • $g(u_{|U_{0}|-1})=m_{k_{1}}$.
If $0<j<j_{0}$, then we note that $f$ maps the interval $M_{j}:=\{m_{k_{j}},m_{k_{j}+1},\ldots,m_{k_{j+1}}\}$ onto an interval containing $f(m_{k_{j}})=l^{-}_{i_{j}-1}$ and $f(m_{k_{j+1}})=l^{-}_{i_{j+1}-1}$, moreover, this interval is a subset of $\{l^{-}_{0},l^{-}_{1},\ldots,l^{-}_{i_{j+1}-1}\}$. On the other hand, $P_{j}\cup\{p_{|P_{<j}|-1}\}$ has endpoints $p_{|P_{<j}|-1}$, $p_{|P_{<j}|+|P_{j}|-1}$, and is mapped by $h$ onto the interval between $l^{-}_{i_{j}-1}=h(p_{|P_{<j}|-1})$ and $l^{-}_{i_{j+1}-1}=h(p_{|P_{<j}|+|P_{j}|-1})$ by $(\bigstar)_{j}$. But $(\bigstar)_{j}$ also says that $l^{-}_{i_{j}-1}$ $l^{-}_{i_{j+1}-1}$ are endpoints of the interval $\{l^{-}_{0},l^{-}_{1},\ldots,l^{-}_{i_{j+1}-1}\}$. So by the above,
$f[M_{j}]=h[P_{j}\cup\{p_{|P_{<j}|-1}\}]=\{l^{-}_{0},l^{-}_{1},\ldots,l^{-}_{i_{j+1}-1}\},$
and both $f$ and $h$ map endpoints to endpoints. So again by Lemma B.2 we define
  1. $(\hexstar)_{j}$
    $U_{j}=\{u_{|U_{<j}|},\ldots,u_{|U_{<j}|+|U_{j}|-1}\}$, and $f^{*}{\restriction}U_{j}\in{\rm Epi}(U_{j},P_{j}\cup\{p_{|P_{<j}|-1}\})$, $g{\restriction}U_{j}\in{\rm Epi}(U_{j},M_{j})$ with
    • $f^{*}(u_{|U_{<j}|})=p_{|P_{<j}|-1}$,
    • $f^{*}(u_{|U_{<j}|+|U_{j}|-1})=p_{|P_{<j}|+|P_{j}|-1}$,
    • $g(u_{|U_{<j}|})=m_{k_{j}}$,
    • $g(u_{|U_{<j}|+|U_{j}|-1})=m_{k_{j+1}}$.
This way
  1. $(\hexstar)_{U}$
    we defined
    • $U=\bigcup_{j<j_{0}}U_{j}$ (as a disjoint union, with $U_{j}$ connected to $U_{j+1}$),
    • the epimorphism $f^{*}{\restriction}U$, which maps $U$ onto $P$, with $f^{*}(u_{0})=p_{0}$, $f^{*}(u_{|U|-1})=p_{|P|-1}$,
    • the epimorphism $g{\restriction}U$ mapping $U$ to the interval $\{m_{0},m_{1},\ldots,m_{k_{j_{0}}}\}$ in $M$, with $g(u_{0})=m_{0}$, $g(u_{|U|-1})=m_{k_{j_{0}}}$.
    Moreover, $h\circ f^{*}{\restriction}U=f\circ g{\restriction}U$, so (by $(\bigstar)_{j_{0}}$)
    (B.4) $f(g(u_{|U|-1})=h(f^{*}(u_{|U|-1})=h(p_{|P|-1})=l^{-}_{|L|-1}(=l^{-}_{i_{j_{0}}-1}).$
  1. $(\hexstar)_{V}$
    We construct $V=\{v_{0},v_{1},\ldots,v_{|V|-1}\}$,
    • $g{\restriction}V$ is a bijection between $V$ and a subinterval of $M$ with $g(v_{0})=m_{-}$ (which is $m_{|M|-1}$ according to our labeling from (B.3)), and $m_{j^{\prime}}$, where $j^{\prime}$ is maximal such that
      $f{\restriction}\{m_{j^{\prime}},m_{j^{\prime}+1},\ldots,m_{|M|-1}\}\textrm{ maps onto }L,$
      equivalently, $f(m_{j^{\prime}})=l^{+}_{|L|-1}$,
    • $f^{*}{\restriction}V\in{\rm Epi}(V,Q)$ such that $h\circ f^{*}{\restriction}V=f\circ g{\restriction}V$, so mapping $V$ onto $Q$ with $h(f^{*}(v_{0}))=l^{+}_{0}$, $h(f^{*}(v_{|V|-1}))=l^{+}_{|L|-1}$, so necessarily $f^{*}(v_{|V|-1})=q_{|L|-1}$ (which can be done in a unique way, as $h{\restriction}Q$ is a bijection).
    Moreover, the fact $h\circ f^{*}{\restriction}V=f\circ g{\restriction}V$ (and $(\nabla_{Q})$$(\nabla_{Q})_{4}$) implies
    (B.5) $f(g(v_{|v|-1})=h(f^{*}(v_{|V|-1})=h(q_{|L|-1})=l^{+}_{|L|-1}$.
It remains to define the linear graph $W$ whose endpoints are connected with $u_{U|-1}$ and $v_{|V|-1}$, respectively, and define $f^{*}{\restriction}W:W\to S$ and $g{\restriction}W$.
$W$ will be of the form
  1. $(\hexstar)_{W}$
    $W=\{w_{0},w_{1},\ldots,w_{|W|-1}\}$ with $w_{i}Rw_{i+1}$,
    • and its endpoint $w_{0}$ connected with $u_{|U|-1}$, while the other endpoint $w_{|W|-1}$ is connected with $v_{|V|-1}$,
    • and we let $f^{*}(w_{0})=s_{0}$, $g(w_{0})=g(u_{|U|-1})$,
    • moreover let $f^{*}(w_{|W|-1})=s_{|S|-1}$, $g(w_{|W|-1})=g(v_{|V|-1})$.
It is easily checked (e.g. by (B.4), (B.5), $(\bigstar)_{S}$) that $f\circ g$ agrees with $h\circ f^{*}$ on the endpoints of $W$ i.e. on $w_{0}$, $w_{|W|-1}$, and the attained values are endpoints of $L$:
$f(g(w_{0}))=h(f^{*}(w_{0})=l^{-}_{|L|-1},$
$f(g(w_{|W|-1}))=h(f^{*}(w_{|W|-1})=l^{+}_{|L|-1}$.
If these two are opposite endpoints, then we can apply Lemma B.2 (to $L$, $h{\restriction}S\in{\rm Epi}(S,L)$ and the subinterval of $M$ between $g(u_{|U|-1})=g(w_{0})$ and $g(v_{|V|-1})=g(w_{|W|-1})$) and so we obtain $W$, $f^{*}{\restriction}W$, $g{\restriction}W$.
If this is not the case, then necessarily $f(g(w_{0}))=f(g(w_{|W|-1}))$, and then pick $m^{\prime}\in M$ with $f(m^{\prime})$ being the other endpoint of $L$. This case we split $S$ into two, $S=S_{0}\cup S_{1}$, where $S_{0}=\{s_{0},s_{1},\ldots,s_{|L|-1}\}$, $S_{1}=\{s_{|L|-1},s_{|L|},\ldots,s_{2|L|-1}\}$. as well as find $W$ in the form of $W_{0}\cup W_{1}$. Note that this case $h(s_{0})=h(s_{|S|-1})$, so (by $(\nabla_{S})$$(\nabla_{S})_{4}$, $(\nabla_{S})$$(\nabla_{S})_{5}$) we have $f(m^{\prime})=h(s_{|L|-1})$. We apply Lemma B.2 twice:
  • to $L$, $h{\restriction}S_{0}\in{\rm Epi}(S_{0},L)$ and $f$ restricted to the interval in $M$ between $g(u_{|U|-1})=g(w_{0})$ and $m^{\prime}$ (to obtain $W_{0}$ and $g{\restriction}W_{0}$, $f^{*}{\restriction}W_{0}$)
  • and to $L$, $h{\restriction}S_{1}\in{\rm Epi}(S_{1},L)$ and $f$ restricted to the interval between $m^{\prime}$ and $g(v_{|V|-1})=g(w_{|W|-1})$, with which we get $W_{1}$, $g{\restriction}W_{1}$, $f^{*}{\restriction}W_{1}$.
It is straightforward to check that $M^{*}=U\cup V\cup W$, $g$, $f^{*}$ are as required (where the endpoints of $M^{*}$ are $m^{*}_{+}=u_{0}$ and $m^{*}_{-}=v_{0}$). ∎
We are ready to prove Lemma B.1.
@@ 1. $(\hexstar)_{W}$
It is straightforward to check that Lemma B.3 together with Lemma B.2 imply the main part of Lemma B.1.

### v2

@@ 2. $f:J\to I$, $f^{\prime}:J^{\prime}\to I$ are epimorphisms (in $\mathcal{P}$}),
So replacing $L$ with the subgraph that remains upon removal of the two endpoints (and letting $g$, $g^{\prime}$ be the appropriate restrictions of $\tilde{g}$, $\tilde{g}^{\prime}$) works. ∎
@@
The proof is by induction on $|I|$ using Lemma B.3 above.
If $|I|=1$, then any $L$ that admits epimorphisms to $J$ and $J^{\prime}$ works. Suppose that Lemma B.2 holds if $|I|<n$, and fix $I$, $J$, $J^{\prime}$, $f$, $f^{\prime}$ where $|I|=n$. Enumerate $I$ as $\{a_{0},a_{1},\ldots,a_{n-1}\}$ such that
$\{\{a_{j}:\ j\leq k\}:k<n\}={\rm tp}^{a_{j},J}(f)={\rm tp}^{a_{J^{\prime}},J^{\prime}}(f^{\prime}),$
(in particular, $f(a_{J})=f^{\prime}(a_{J^{\prime}})=a_{0}$), and pick an enumeration $\{b_{0},b_{1},\ldots,b_{n-1}\}$ with
$\{\{b_{j}:\ j\leq k\}:k<n\}={\rm tp}^{b_{J},J}(f)={\rm tp}^{b_{J^{\prime}},J^{\prime}}(f^{\prime}).$
We note that both $a_{n-1}$ and $b_{n-1}$ must be (possibly different endpoints) of $I$. We assume that $a_{n-1}=b_{n-1}$ (the other case is similar but simpler, and is left to the reader).
Now we reduce the task of finding $I$ and $g$, $g^{\prime}$ to four amalgamation task. Enumerate
$J=\{a_{J}=u_{0},u_{1},\ldots,u_{|J|-1}=b_{J}\}\ \textrm{ so that }u_{k}Ru_{k+1},$
and
$J^{\prime}=\{a_{J^{\prime}}=u^{\prime}_{0},u^{\prime}_{1},\ldots,u^{\prime}_{|J^{\prime}|-1}=b_{J^{\prime}}\}\ \textrm{ so that }u^{\prime}_{k}Ru^{\prime}_{k+1}.$
Let
  • $s_{0}<|J|$ be minimal such that $f(u_{s_{0}})=a_{n-1}$, and
  • $s^{\prime}_{0}<|J^{\prime}|$ be minimal such that $f^{\prime}(u^{\prime}_{s_{0}^{\prime}})=a_{n-1}$,
and set
$K_{0}=\{u_{0},u_{1},\ldots,u_{s_{0}-1}\},$
$K_{0}^{\prime}=\{u^{\prime}_{0},u^{\prime}_{1},\ldots,u^{\prime}_{s^{\prime}_{0}-1}\}.$
While it is clear that
(B.3) $f[K_{0}]=f^{\prime}[K_{0}^{\prime}]=I\setminus\{a_{n-1}\},$
(B.4) ${\rm tp}^{u_{0},K_{0}}(f)={\rm tp}^{u^{\prime}_{0},K_{0}^{\prime}}(f^{\prime}),$
we argue that
(B.5) ${\rm tp}^{u_{s_{0}-1},K_{0}}(f)={\rm tp}^{u^{\prime}_{s^{\prime}_{0}-1},K_{0}^{\prime}}(f^{\prime}),$
holds as well. Indeed, by the way we defined $s_{0}$ and $s^{\prime}_{0}$ we have $f(u_{s_{0}-1})=f^{\prime}(u^{\prime}_{s_{0}^{\prime}-1})=a_{n-1}$ is and endpoint of $f[K_{0}]=f^{\prime}[K_{0}^{\prime}]=I\setminus\{a_{n-1}\}$, so it is enough to recall that there is only one type starting with a fixed endpoint.
Similarly, we let
  • $t_{0}<|J|$ be maximal such that $f(u_{t_{0}})=b_{n-1}$, and
  • $t^{\prime}_{0}<|J^{\prime}|$ be maximal such that $f^{\prime}(u^{\prime}_{t_{0}^{\prime}})=b_{n-1}$,
and set
$K_{3}=\{u_{t_{0}+1},u_{t_{0}+2},\ldots,u_{|J|-1}\},$
$K_{3}^{\prime}=\{u^{\prime}_{t^{\prime}_{0}+1},u^{\prime}_{t^{\prime}_{0}+2},\ldots,u^{\prime}_{|J^{\prime}|-1}\}.$
So again,
(B.6) $f[K_{3}]=f^{\prime}[K_{3}^{\prime}]=I\setminus\{b_{n-1}\},$
(B.7) ${\rm tp}^{u_{|J|-1},K_{3}}(f)={\rm tp}^{u^{\prime}_{|J^{\prime}|-1},K_{3}^{\prime}}(f^{\prime}),$
(B.8) ${\rm tp}^{u_{t_{0}+1},K_{3}}(f)={\rm tp}^{u^{\prime}_{t^{\prime}_{0}+1},K_{3}^{\prime}}(f^{\prime}).$
Now using B.3-B.5 and applying the induction hypothesis
  1. $({\rm R}0)$
    $f{\restriction}K_{0}$, $f{\restriction}K_{0}^{\prime}$ can be amalgamated in a way that endpoints are mapped to endpoints, more concretely one endpoint is mapped to $u_{0}$ and $u^{\prime}_{0}$, while the other is mapped to $u_{s_{0}-1}$ and $u^{\prime}_{s^{\prime}_{0}-1}$,
and by (B.6)-(B.8),
  1. $({\rm R}3)$
    $f{\restriction}K_{3}$, $f{\restriction}K_{3}^{\prime}$ can be amalgamated in a way that endpoints are mapped to endpoints, more concretely one endpoint is mapped to $u_{|J|-1}$ and $u^{\prime}_{|J^{\prime}|-1}$, while the other is mapped to $u_{t_{0}+1}$ and $u^{\prime}_{t^{\prime}_{0}+1}$.
Recall that we assumed that $a_{n-1}=b_{n-1}$ (which is an endpoint of $I$), pick $q_{0}<|J|$ and $q^{\prime}_{0}<|J^{\prime}|$ such that
(B.9) $f(u_{q_{0}})=f^{\prime}(u^{\prime}_{q_{0}^{\prime}})\textrm{ is the other endpoint.}$
Set
$K_{1}=[u_{s_{0}},u_{q_{0}}],$
$K_{1}^{\prime}=[u^{\prime}_{s^{\prime}_{0}},u^{\prime}_{q^{\prime}_{0}}],$
and
$K_{2}=[u_{q_{0}},u_{t_{0}}],$
$K_{2}^{\prime}=[u^{\prime}_{q^{\prime}_{0}},u^{\prime}_{t^{\prime}_{0}}].$
So by the way $s_{0}$ and $s_{0}^{\prime}$ are picked $f(u_{s_{0}})=f^{\prime}(u^{\prime}_{s^{\prime}_{0}})$ we note that $f{\restriction}K_{1}$ and $f^{\prime}_{\restriction}K_{1}^{\prime}$ map endpoints to endpoints, so Lemma B.3 applies, so
  1. $({\rm R}1)$
    $f{\restriction}K_{1}$, $f{\restriction}K_{1}^{\prime}$ can be amalgamated in a way that endpoints are mapped to endpoints, more concretely one endpoint is mapped to $u_{q_{0}}$ and $u^{\prime}_{q_{0}^{\prime}}$, while the other is mapped to $u_{s_{0}-1}$ and $u^{\prime}_{s^{\prime}_{0}-1}$.
Similarly,
  1. $({\rm R}2)$
    $f{\restriction}K_{2}$, $f{\restriction}K_{2}^{\prime}$ can be amalgamated such that endpoints are mapped to endpoints, more concretely one endpoint is mapped to $u_{q_{0}}$ and $u^{\prime}_{q_{0}^{\prime}}$, while the other is mapped to $u_{t_{0}+1}$ and $u^{\prime}_{t^{\prime}_{0}+1}$.
Putting together the maps in $({\rm R}0)$, $({\rm R}1)$, $({\rm R}2)$, $({\rm R}3$ we get the conclusion. ∎
Proof of Lemma B.1.
We reduce Lemma B.1 to Lemma B.2.